In experiment 2, we will be measuring the time it takes for a marble to roll different distances. We will plot the distance the marble moves LaTeX: d d as a function of LaTeX: t^2 t 2 . If the plot is a line (or a fair approximation), this means that we have constant acceleration. If the plot of LaTeX: d d as a function of LaTeX: t^2 t 2 has a slope of 2.73 m/s2, what is the acceleration? Please give your answer in units of m/s2 and 2 decimal precision.
Answer #1

d = at^2 / 2

Where a is acceleration of marble.

In a plot of d as function of t^2, slope of line is a/2. Hence

Slope 2.73 m/s^2 = a/2

a = 5.46 m/s^2